OpenStudy (anonymous):

Write the polynomial in factored form. 4x^3 + 8x^2- 96x

OpenStudy (jazzyluvsu123):

Explanation: 4x^3, 8x^2 and 96x have 4x as a common factor. Hence 4x^3+8x^2-96x = 4x*(x^2+2x-24) Furthermore, (x^2+2x-24)= (x-4)(x+6) because 6-4=2 and 6*(-4) = -24. So 4x^3+8x^2-96x = 4x*(x^2+2x-24) = 4x(x-4)(x+6).

OpenStudy (anonymous):

Thank you so much @JazzyLuvsU123

OpenStudy (jazzyluvsu123):

Your very welcome @Madelynrawls